OpenStudy (anonymous):

Please find the general solutions to 6cos^2(2x)-3=0

OpenStudy (anonymous):

\[2\cos ^2x=\frac{ 3 }{ 6 }*2=1,1+\cos 4x=1,\cos 4x=0\] can you find the angle where cos4x=0

OpenStudy (anonymous):

correction\[2\cos ^2(2x)\]

OpenStudy (anonymous):

it is on y axis.

OpenStudy (anonymous):

cos^2(2x)=1/2

OpenStudy (anonymous):

\[\cos 4x=\cos \frac{ \left( 2n+1 \right)\pi }{ 2 },4x=\left( 2n+1 \right)\frac{ \pi }{ 2 },x=?\] here n is an integer.

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you able to find the solution?

OpenStudy (anonymous):

no i made it to cos^2(2x)=1/2

OpenStudy (anonymous):

do you square the 1/2 to get (√2/2)

OpenStudy (anonymous):

then to 2x=pi/4+2npi

OpenStudy (anonymous):

\[x=\frac{( 2n+1)\pi }{ 8 }\]

OpenStudy (anonymous):

\[\cos 2x=\pm \frac{ 1 }{ \sqrt{2} }\] to avoid negative sign i have used double angle formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

yw.